Cavity alignment was not recoverable. I most likely spent way too much time trying to recover instead of just starting from scratch. This is frustrating for future first contact cleaning if this cleaning procedure does improve the cavity finesse. The proccess I tried was the following:
1. Assume input alignment was unchanged, place both input couplers back on the cavity. Adjust input couplers to align both light paths back on their refl PDs. With how small the diode is on the high BW REFL PDs this should ensure 2 mirrors in their original position.
2. Piezo mirror (775 output coupler) is static and cannot be changed. Nothing to be done there.
3. Scan around on third adjustable mirror and look for any second pass beams on the camera.
It seems to me like this should work, but this did not yield anything.I believe the piezo mirror is at a drastically different angle, which is strange because when I was doing piezo science before, changing out the piezos did not drastically change the intercavity alignment. I started over by changing the input alignment of the 1550 light and adjusting the cavity mirrors to match the 3D printed alignment tools. Doing this instead I found cavity flashes within an hour. Realigned the REFL and TRANS light onto the cameras/PDs and fine tuned alignment. OFC1 1550 light is now aligned. With the newly built tank circuit the error signal on the 1550 path looks very clean. I believe the frequency of the modulations is not quite correct and can be looked into further. I have taken ringdown measurements to see if cleaning was successful, but am getting finesse numbers greater than the specs of the mirrors. I think this may be due to the TRANS pd being saturated? Going to retry. Also need to double check the 1550 TRANS PD is fast enough, although for the 1550 light the decay time should be on the order of 3300/pi/c*2.4 = 8e-6 s or 100's of kHz. I think all the PDs we have should be this fast.
Additionally, i took off the 775 input coupler and tried to match the input alignment. As built, the 775 light first sees the piezo mirror after the input coupler. For any given input alignment the piezo mirror is putting the beam into the metal casing of the cavity. Need to explore this still.
Super optic cleaning prep notes:
I've moved the bin of extra OFC parts to the other table in order to clear room on the OFC table. This will allow us to keep the exposed super optics inside the clean room at all times. I've set up a prep area on the OFC table closest to the door to B102. I have cleaned a pair of metal tweezers (the plastic ones in trials were very staticy), all of the required hex keys, and moved the first contact supplies to this location.
Currently my plan is to fully remove the ~2 inch spacer/flexture mount combination along with the optic. It would be possible to unscrew the nylon tip set screw that holds the mirrors in place but this seems like it adds unneeded risk and the flexture mount combinations can be removed/replaced and still retain inter-cavity alignment.
The piezo mirror must be fully removed in order to access the mirror. I have set aside an SM1 ring spanner wrench and piezo alignment tool to be cleaned when I am garbed up.
Two of the mirrors on one side (OFC1:FCM2 and OFC1:FCM4) are too close to a beamsplitter (OFC1:BS2) to remove. This BS must be removed from the table temporarily. I have done this.
End of preparation for first contacting particle count 0/0/0 and 1/0/0 during two 60 second count trials.
** Lunch**
Midway particle count: 0/0/0
Plan is to clean both the HR and AR sides. I think the way to do it is rest the optic on the AR side, paint the HR side, let this side dry, flip it and paint the AR side while the optic is resting on the painted HR side, wait til both are dry, remove both sides without resting the optic and install back into mount.
HR side first contact complete. Piezo mirror was visibly dirty. Looked like metal shavings. The AR side also had a large imprint on it from the piezo being driven a couple of times. Similar to what we've seen in the past.
AR side first contact complete. Mirrors are resting on dry HR side. Waiting 10 more minutes and then peeling.
Cleaned the inside circular part of the spacers that hold the mirror. The piezo mirror mount one had what looked like small metalic particulates.
All 8 peels were very successful. I noticed some specs (most likely stringing from the FC) on the AR side of OFC1:FCM1. Repainted this side while it was in its mount. Peel was successful. Reinstalled the mirrors on the cavity and now will scan and look to realign.
End of work particle count: 0/0/0
[Jeff, Daniel, Sander]
We have moved the TMC optical tables into close to their 'final' positions, according to the designed lab layout (see attached). We extended the caster wheels in the tables, which allows them to be moved around with two people pushing and one person guiding the table. The east table is not pushed up to the other table fully. We need to find some way to do this without bumping the tables into eachother and end up with both tables in the final desired positions. Note that the table tops are raised when the wheels are extended (vs. wheels retracted) and that the breadbord table tops protrude slightly from the base of the top (i.e. there's a horizontal 'lip' at the top), which makes this more difficult.
Some observations:
- As mentioned previously, the tabletops are not rigidly attached to the legs. This means the tops can and will move slightly on the tops of the pistons (which are on top of the legs) that the tabletop rests on, limited only by friction. The only thing preventing the top from moving completely off the legs are the earthquake stop bars.
- It seems the heads of the pistons (black part in photo), are not aligned with the bodies of the pistons (silver part in photo); inadvertently moving the top w.r.t the legs probably caused this misalignment. This misalignment is quite bad for some of the legs (see photo).
- The table tops do not rest equally on all four legs, there are visible gaps between the tops of the pistons and the top (see photo). This is probably because the lab floor is not level.
I think we can do the final movement by extending wheels on both tables, so that the breadbord lips are at the same height and the tables can recoil when pushing them together. Then we can try to slowly push the tables together so there's no gap, which will likely move the tables away from their desired final position. Then we could put e.g. a breadbord over the seam to make a rigid connection, after which we can do the final positioning, moving both tables at the same time.
You have the SHG sled and power distribution on this diagram twice just fyi.
[Alex, Torrey, Daniel]
We made and tested the amplifiers for the GQuEST AOMs. They work well.
Steps:
One end should be unused
Shielded BNC solder procedure:
Use 10 gauge solid, go 3/4” down, strip plastic away
Remove some extra ground strand
Twist exposed ground strand together
Use 14 gauge stripper, expose ~3/8” of copper wire
When soldering, make sure only pin touches BNC
I still want to add some wire mesh to shield the back from RF but allow for air to flow to the amplifiers.
Alex made the power supply seperate from this chassis. I will let him comment on how to make that.
I've set up an account number for techmart for our lab to purchase high purity N2 from AirGas. The account number is 4890116.
[Briana, Ian, Torrey]
Probe measurements again:
Retook spectra with just the probe to compare the results from using the matplotlib psd function (will switch to numpy), shown here probe_psd_python.png, and the Moku spectrum analyzer, shown here: probe_psd_moku.png. Data was taken for the photodetector input Moku channel at 50 Ohm impedance and 1 MOhm impedance to test at "different gains," although this does not really affect digitization noise since the signal would have already been digitized upon reaching the Moku- changing gain would have to happen before the Moku entirely. Regardless, the units of the y-axis are different between these two images but the order of which has the highest to lowest noise should be the same. However, the Moku spectrum analyzer provides less resolution so the shapes (especially <10 Hz) are not easily comparable and I am generally having a hard time trusting the spectrum analyzer. The general order of highest to lowest noise is somewhat consistent at <10 Hz. For 50 Ohms, after moving the offset of the temperature controller completely off the temperature at which the absorption occurs, we somehow see less noise compared to when the laser is locked, which is very bizarre. It is possible that the laser may just become very stable at that random temperature that I detuned to. Also, the noise level for the spectrum analyzer at higher frequencies is offset in the spectrum analyzer, in contrast to the post-processing method which shows what we would expect: noise at higher frequencies is the same across all three cases because the actuator does not implement control at those frequencies. I'm not sure why the spectrum analyzer behaves like this.
Attempts to amplify signal: We want to amplify the signal to reduce effects of bit noise, even though we amplify other noise in the process. Dominating bit noise could be one cause of these confusing noise spectra.
Built tank circuit for the EOM to enhance the error signal as much as possible. Resonant EOM seems to occur at 123 MHz after optimizing phase shift for different modulation frequencies. The EOM has capacitance of 14 picofarads. For a resonant frequency of 122.9 MHz, this means we get an inductance of L = (1/2pi f)^2 / C ~ 119 nano Henries. The resulting tank circuit did not work: touching the BNC cables even slightly changed the error signal significantly, and the signal was not amplified- in fact, the amplitude was larger without the tank circuit. We will proceed without the tank circuit for current time constraints. Connections are fine in the circuit (tested with the silver remote box), but maybe the components/connections are not secure and thus extremely sensitive to the environment/movement.
We want to amplify the light passing through the vapor cell before reaching the Moku (i.e. getting digitized), so we tried adding a low noise amplifier (LNA). I stole the one from the new lab which somehow works better (less noisy signal) than the one in the beige cabinet in B102 even at the same settings. However, amplifying the signal (with lowpass filter on the LNA set to the highest setting of 1 MHz) caused the error signal to become <20 microvolts in amplitude compared to the previous >400 microvolts in amplitude without the LNA, no matter how the modulation frequency/phase is tuned. The error signal also becomes a distorted symmetric shape. I think this is a filtering issue with the lowpassing. Although the photodetector signal from the temperature scan looks the same as without the LNA, maybe the modulation signal strength is not strong enough (the one getting mixed with the photodetector signal) to produce a significant amplitude for the error signal. Additionally, the error signal is very sensitive to moving the BNC cable to the EOM when connected to the LNA. Without the LNA, the error signal is not sensitive to BNC cable movement, so this is some issue with the LNA. Some of its specifications (from here) should also be considered when troubleshooting, although I don't think they are the sole reason for these problems: 1). the output type is a 100 Ohm BNC, so we might want a 100 Ohm BNC cable attached instead of a 50 Ohm (?) and 2). the input to the LNA must be between -0.4 and 0.4 V to get accurate readings (this is fine with our input signal, which usually doesn't even reach 200 mV). We are proceeding without the LNA for now. One idea to troubleshoot would be forking two outputs from the LNA into the Moku, one AC and one DC coupled. AC removes the DC offset so this coupling means that only AC signals will pass through, effectively acting as a lowpass filter. If the error signal is prominent in AC coupling, it could mean we just have an issue with the settings for modulation frequency/phase/low-pass filter in the Moku.
Pump:
Took data of the pump, will analyze this to see what hyperfine transitions they correspond to. Pump fell out of alignment so needs realigning. On Tuesday, we achieved lock! We took 20 seconds of 5000 samples per second for the locked and unlocked cases using the same method as with the probe and we get the following PSD: pump_noise_spectrum.png. The unlocked case has higher noise than the unlocked case, which is what we would expect. It is not immediately obvious why this seems to behave as we expect and not the probe, but I wouldn't trust it yet since we did this very quickly and only once so far (also the modulation frequency/phase/alignment were not optimized). Slow controller used and error signal: controller_20240813_193716_Screenshot.png. Blue dot in image shows the peak that we locked to. Also, the standard deviation of the error signal when locked was ~2.12 e-6 V whereas when unlocked, it is 3.72 e-6 V, which also matches with what we expect since the error signal would be greater and fluctuate more for an uncontrolled unlocked laser.
Current setup: IMG_2454.jpg
[Torrey, Ian]
A theory on the poor finesse measured on the OFCs has been that the mirrors are simply dirty. After hearing Hartmut mention that their optics were dirty straight out of the box from the supplier has convinced me to move forward on cleaning the OFC super optics. Koji has kindly agreed to oversee on Monday at 1 PM. In preperation for this I am going to take up to date measurements of the filter cavities to directly compare to after the cleaning. Currently the plan is to take an updated mode scan, ringdown measurement, and power budget (this has some flaws due to the poor lock quality but we will do our best). In addition, a quick inventory check before this happens:
-Fresh First contact - Yes
-Peek Mesh - Yes
-Clean scissors - Can clean a pair ahead of time
-Clean Forcepts/Tweezers - Same as above
-High purity N2 with a regulator - this we do not have. I think this can be purchased on campus and obtained very quickly but would ask Lee's approval.
-Top Gun - Yes.
In terms of data taking:
Cav Scan and updated ringdown measurements can be seen in "Nextcloud\GQuEST\B102\Output Filter Cavities\PreFirstContactData\"
OFC1
-1550 input - 15.6 mW
-1550 trans - 3.7 mW * 2 (through 50:50 BS)
-1550 refl - 5.2mW
I do not think the 1550 path loss is this bad. As previously discussed it is hard to optimize the 1550 output while locked on 775.
-775 input - 50 uW
-775 REFL - 4 +/-1 uW
-775 TRANS 46 uW
Update to this. Koji came into the lab and gave me some useful tips for first contacting. I've also been emailing with Koji and Garilynn. Couple points for everyone:
-AirGas (the vendor) is on techmart, this is who people use to buy high purity N2 for first contact purposes. I've set up an account for the Lee Lab. Should be able to see it on the website later today.
-Conversly Garilynn thinks its probably fine to clean these small mirrors without N2. The point of N2 is when peeling the dried first contact it can cause static build up, this is not as big of a problem with the small mirrors we are using.
-Last point that is worth mentioning is that LIGO "cleans and bakes" the mesh they use to peel FC. Garilynn pointed out an alternate peel method of putting a small drop on the edge of the optic and use this to initiate the peel.
We are currently wondering why the 1550 nm light is not co-resonant with the 775 nm light, used to lock the GQuEST Output Filter Cavities, over time. Some ideas are the AOM frequency drifts (unlikely), but a potential culperate is differential length changes of the cavity due to thermooptic effects (thermal expansion and thermo-refraction). For light that hits the surface of the coating, Evans (2008) (Eq. 3) gives the sensitivity of the sensed position of the mirror \Delta T to temperature fluctuations T as
\[ \frac{\partial \Delta z}{\partial T} = \bar{\alpha}_c h_c - \bar{\beta}\lambda \]
Where \bar{\alpha}_c is the effective coefficient of thermal expansion, h_c is the coating thickness, \bar{\beta} is the effective thermo-refractive coefficient, and \lambda is the wavelength. This expresion doesn't work for the 775 nm light since it is below a layer of 1550 nm coating. Assume a fraction (\gamma) of the coating is for 1550 nm light and (1-\gamma) of the coating is for 775 nm light. \gamma \approx 0.5, although the 775 nm light coating is probably thinner since the fractional wavelength stacks are thinner. This analysis might also not be exact for the mirrors since none of them are HR for both wavelengths.
\[ \frac{\partial \Delta z_{775~\text{nm}}}{\partial T} = (1-\gamma)\bar{\alpha}_c h_c - \gamma\bar{\alpha}_c h_c(n-1) - \bar{\beta}( \gamma h_c+\lambda) \]
Where n is the index of refraction of 775 nm light in the 1550 nm coating. There is a minus sign in front of it because a higher index causes more phase to be accumulated than in air.
Now consider the differential sensitivity of the sensed position of the mirror, assuming the thermorefractive coefficient is very roughly the same for both wavelengths.
\[ \Delta \frac{\partial \Delta z}{\partial T} \equiv \frac{\partial \Delta z_{1550~\text{nm}}}{\partial T} - \frac{\partial \Delta z_{775~\text{nm}}}{\partial T}\]
\[ \Delta \frac{\partial \Delta z}{\partial T} = \gamma h_c (\bar{\alpha}_c n + \bar{\beta}) \]
Plugging in numbers from our GQuEST Paper Table II,
\[ \Delta \frac{\partial \Delta z}{\partial T} \approx 2 \cdot 10^{-10}~\text{m/K} \]
I'm not sure if ~1/1000 of a wavelength can explain the drift in the co-resonance over time.
Using Delta f /f = Delta l / L, where f is the frequency of the light and L is the length of the cavity, we get 16 kHz / K, not enough to explain the overnight drift of ~ 1 MHz. Lee points out that this temperature dependence might be enhanced by the quarter and half wavelength stacks.
-There is no beat note if the tank circuit on the 1550 EOM is removed.
-If the tank circuit for the 775 is removed the same 120 Hz oscillations are observed. The PDH error amplitude is also very small, as expected.
-I am going to remake a tank circuit and plan to operate OFC1 at 70 MHz (~20 MHz difference, should be filtered out by the low pass in the laser lock box).
Maybe it's fine, but a 20 MHz beat notes seems dangerously close to our nominal readout frequency. However, this is still ~100x the single cavity pole so it should be filtered plenty.
Torrey asks here whether the Output Filter Cavity Offset Frequency can fluctuate from the Nominal Value of 17.6 MHz. If we assume 4 cavities have the same single cavity bandwidth of \Delta epsilon_1 and the same frequency offset, then the integrated transmission is around \Delta epsilon_1 / 2 (see section A.9 in our GQuEST paper). If we assume that 3 cavities have the same frequency offset and assume an allowable additional 5% signal transmission loss due to this mechanism, then the frequency can shift 16% of the bandwidth away from the nominal offset.
I derived this by integrating 4 Lorentizans, once with all 4 with the same offset and once with one with a deviated offset. I then figure out the deviation required so that the transmitted bandwidth is 95% of the no deviation case. For our bandwidth of 42 kHz, this is a value of 6.6 kHz.
If this single cavity had a fluctuating offset, then the RMS discrepancy between this cavity could probably be \sqrt{2} larger because it spends time near no offset deviation. I am assuming there are no dynamic effects of changing the offset when the round trip time in the cavity times the frequency fluctuation is much less than 1, which is the case becuase t_cav*f_bandwidth = 1/2pi.
If I make an assumption that at any given time the 4 cavities are randomly distributed, let's say "equally spaced" on a gaussian with z-scores of -0.84, -0.25, 0.25, and 0.84 (the integral from one value to the next, including +/- infinity, is 0.2), then the 1 sigma error for 95% transmission is 11% of the bandwidth away from the nominal offset, or 4.6 kHz. This is only a rough estimate and more sophisticated methods should be used if we want a more exact answer, but a good sanity check is that it is below the 1 cavity answer.
One should note for this model that I'm assuming some delta f_RMS since this is a noise issue, not a static frequency alignment issue. In terms of converting this frequency to a length requirement, we use
\[ \frac{L_{\text{cav}}}{f_{\text{Laser}}} = \frac{l_{\text{RMS}}}{f_{\text{RMS}}} \]
\[ l_{\text{RMS}} = f_{\text{RMS}} \frac{L_{\text{cav}}\lambda_{\text{Laser}}}{c} \]
L_cav = 2.4 m, \lambda_Laser = 1550*10^-9 m, and c is the speed of light. Plugging in f_RMS = ~5 kHz, we get l_RMS = 6*10^-11 m = 0.06 nm = 0.6 Å. This is an order of magnitude smaller than wavelength over the Finesse of the cavity (F = 3000), which is 0.5 nm = 5 Å.
For OFC2, the laser lock quality seems poor. This and this show the 775 error signal used to lock the cavity in red, and blue shows the TRANS PD for 1550. I am not sure how to quantify a "good" lock in these cavities. For any diagnostics, this is a poor lock quality as the total 1550 transmission is fluctuating by alot. But in GQuest operation mode we will have it detuned by 17.6 MHz, thus supressing 1550 light. If it jitters around 17.6 MHz, still suppressing the carrier and allowing signal photons through, is this still an issue? We cannot suppress those fluctuations further, even though increasing the UGF of the loop squashes these fluctuations. At a certain point, something rings up (in the laser dc modulation port? elsewhere?) and doesn't allow further supression.
I am attempting to look at this issue on OFC1. It is largely the same. One major problem I diagnosed is having both EOM phase modulations on in one cavity at the same time. The clock sync that we have between the 4 mokus seems to not be sufficient. The signal is fluctuating multiple times a second. One moment and the next. This may not seem like a big deal but since the two wavelengths are tracking eachother, having a 1mV difference in the average level of the error point has the effect of detuning the 775 light away from the 1550 peak about a FWHM distance (read: when the average error point is 0, they are coaligned, when the average is lower, the 1550 TRANS light jumps to about half the value). Turning off the 1550 EOM signal gets rid of it. This is when the two EOMs are at the same frequency. When they aren't there is a large beat note in the error signal.
I think we can tolerate a few kHz of frequency jitter. See this.
I would like to do some analysis on the filter cavities but am getting hung up on the fact the 1550 PDH error amplitude is very low. I've double checked the EOM frequency, inductor tuning in the tank circuit, and phase of the PDH loop. All seem fine. I believe it's somehow related to alignment, the REFL dips on the OFC1 1550 path are the lowest out of the 3 REFL PDs I have installed so far ( OFC1 1550, OFC1 775, OFC2 775). Below are screenshots of the error amplitude. If anyone has any other ideas to improve this please let me know. It wasn't this small in the past.
I scanned the laser on OFC2 for the first time since yesterday. There were higher order mode peaks on the 775 transmission PD at values higher than previously observed. This is most likely due to the earthquake that occured very near campus. I touched up the input alignment and now have almost no higher order modes seen on the TRANS PD.
It's most likely 1 optic poorly clamped down that caused this. Will investigate.
[Alex, Daniel]
I loaned out our RFSoc to the Hutzler Lab. The point of contact is Harish Ramachandran. Also in the loan was the included power cord and some SMA connectors.
[Briana, Ian]
7/18: Started taking data with a 150 mm lens in but messed up wavelength in JamMt and alignment so data is void. However, from the initial look at that data, it seemed like putting in the lens removed the finer dips within larger dips (some of the absorption features are getting lost). However, wouldn't trust this too much because the alignment was bad. When I removed the lens, there was a dip in transmittance (blue circle) but by shifting the photodetector around, you can get back to a level close to the original with-lens measurement, although there is some noise (red circle), (see rip_alignment). This leads me to also think the issue in the last post could also be an alignment problem. However, I am not sure as I don't think I touched anything that would move it out of alignment, but who knows.
7/19: Realigned and measured for the 150 mm lens. First, I scanned across temperature with the lens in place. Then I removed the lens and scanned. Then I put the lens back in its original place, beam profiled the vapor cell, and moved the photodetector to see if I could get the same power level. I measured the beam profile on the active area of the photodetector, right before entering the vapor cell, and right after the vapor cell. The issue was the order in which I did this because it required a lot of putting the lens back to where it was, which might not have been accurate. A more consistent procedure is discussed below. At this point I also put in a 0.2 OD ND filter to not saturate the blue nanoscan beam profiler (didn't find a setting to increase saturation limit in the program).
7/21-7/22: Took data with the 100 mm lens (at 25 Celsius on 7/21, at 40 Celsius on 7/22). This is the procedure going forwards (can be repeated for the different focal lengths):
First, profile the beam. The laser power should not affect the beam profile but for consistency, set the laser current at 110 mA, temperature at 6.755 kOhms. Set the position of the lens and vapor cell (everything is relative to the mirror before the beam enters the vapor cell). Record the positions of where the vapor cell will begin/end, where the lens will be. Beam profile at the beginning of the vapor cell without the lens in place and record the location. Put the vapor cell in place. Beam profile after the vapor cell without the lens and record the location. Determine the location of the photodetector where the waist should be. Ensure that the beam profile you measure at the photodetector is within the 3.6 mm by 3.6 mm aperture and record the location/beam profile. Put photodetector down and mark the location. Slightly tune the positioning of the photodetector until you can see good dips (this was largely unsuccessful, I'm not sure how the tuning works). Do the scans/measurements (NO LENS). Put the lens in place. Confirm that the light falling onto the photodetector is the same as without the lens by checking alignment. Do scans (WITH LENS). Put in the beam profiler where the photodetector is (keeping the lens in place), which will require removing the photodetector. Then, beam profile after the vapor cell. Remove vapor cell and beam profile right before the lens.
Using Torrey's beam profiling code and Jammt, you should be able to determine the beam size reasonably accurately. I was running into issues where the beam size I measured differed by >3 mm from JamMt because I should have picked 'thick lens' on the JamMt settings (not thin lens). With this fixed, there is a ~1mm difference in beam size than expected. Could be explained by error in distance measurements. Moving forward, I'll verify the distance measurements more accurately to see if this is the issue. Because I was having trouble with JamMt, I had started profiling the beam in the sequential manner, so I think the values for the beam size are correct, but just are not verifiable with JamMT.
Results:
The dip depth comparisons from all the data are shown here with the measured beam sizes (DipDepthComparison.png). Even at a different vapor cell temperature, the smaller beam size decreases the strength of the absorption dip (the 150 mm lens produces a smaller beam size by about ~1 mm). I want to justify this by saying that a smaller beam size means photons will interact with a lower number of atoms so less photon absorption occurs (in contrast, a larger beam means photons are more likely to be absorbed because they interact with more atoms). I feel like you would be saturating the atoms by bombarding them with photons, but I'm not sure if there are too many atoms for this to occur (spontaneous emission process is too fast?). Something I'll look into/calculate.
Also, if you plot the relative dip depth (1 - background voltage / signal voltage), you see that it decreases as you increase power, which would make sense as the increase in absolute dip depth is not enough to overcome the larger increase in background power (relative_dip_depth). This could also explain the saturation effect that we see (dip depth becomes less pronounced once you increase power too much).
From the 7/19 plots, it seems like there is a power limit before the dip depth begins to decrease, which is due to some saturation effect. I want to say that this is a saturation limit on the photodetector, but as Ian pointed out it could be an issue with the Moku, so I should measure the power before reaching the photodetector to confirm. Based on the manual, if the photodetector is reading over 5V for a 50 Ohm load (which is what the Moku is set to), then it will saturate, but the output has consistently exceeded 5V output reading and has still retained an increasing dip depth. Should double check settings.
Smaller absorption features (on both dips) are still getting lost. If you get the right alignment, you can see the dips. Maybe the photodetector needs to be a certain distance away to register it but I'm not sure why/how since all the light is getting concentrated onto the active area anyways.
The way I measure the dip depth is to take the ends of the temperature vs. reading plot and fit an exponential to it (because the shape of the temperature curve sometimes trails up towards higher temperatures and is not a linear background). After subtracting it, I get something like the second slide. subtracting_background_example.pdf. Before I was doing it by hardcoding indices, so this should be better and easier to verify.
See picture of the setup for the measurements. I plan to do the same with a 75 mm lens for a more drastic difference in beam size. Before that, the laser output will be collimated using the fiber collimators at the point of the alser. Ideally, light exiting the laser would be collimated. This makes it easier to mode match the beam, which we will need to do eventually for the probe/pump.
Miscellaneous:
Connected CTL OUT of TED200C (laser temperature control) to input 3 of Moku (labelled Paige Bueckers). Must set the CTL OUT impedance of at least 10 kOhms from the TED200C manual, otherwise it will not register so on the Moku an impedance of 1 MOhm was set for the channel of CTL OUT input. This output is a voltage proportional to the actual temperature. The voltage range is 0 to 10 V with the conversion coefficient of 2 kOhms/V (using TH 20 K sensor). Using this signal instead of the previous triangular wave signal, which is more reflective of the actual temperature, did not solve the shifting issue: the frequency still has to be low to produce a not obvious shift. I thought this could be because of cable length differences between the CTL OUT BNC cable and the photodetector BNC cable, but probably not since electricity travels extremely fast. I also think that the temperature controller has asymmetry since cooling the laser may not have the same efficiency as heating it (this could be something with the PID controller of the laser temp controller).
I saw this page and thought it was interesting how a hot spot on the laser looks kind of like an inverted absorption dip. I didn't know lasers could do this- it probably shows up in the beam profile.
Wow, that's awesome to hear. Good optics and good procedures.
Update:
Retook the measurement, making sure to move the laser frequency away from equilibrium enough via the output offset, so that when I break lock no new light is resonant and I am purely measuring the light exiting the cavity. Additionally reduced input power slightly so that the PD is not saturated. Did two measurements in transmission of the cavity and find finesse measurements very close to the specs of the mirror. I think more analysis is needed (align 775 and do similar measurements) but we may be able to conclude that the mirrors were simply dirty coming from the manufacturer. Plot attached. It seems like this may be worth doing for the second cavity.